Solve by elimination
2x + 3y =-4
3x-4y=11

Answer by sachi(548)

2x + 3y =-4 ........1
3x-4y=11 ........2
now multiplying 4 with eqn 1 & 3 with eqn 2 we get
8x+12y=-16.............3
9x-12y=33 .............4 now adding both eqns 3 & 4
17x=17
or x=1
from eqn 1
2*1+3y=-4
or y=-2